OpenStudy (anonymous):

6ab^-3/6^-1a^-4b^2, I can't recall the rule for negative exponents with this

OpenStudy (anonymous):

\[x^{-n}=\frac{1}{x^n}\]Does that help?

OpenStudy (anonymous):

I don't see how that applies exactly though. Would I have a 3 tiered fraction??

OpenStudy (anonymous):

It's difficult for me to help directly since the expression isn't clear. Is everything to the right of the '/' under everything to the left?

OpenStudy (anonymous):

I now see the equation box at the bottom lol, one sec

OpenStudy (anonymous):

\[6ab^-3/6^-1a^-4b^2\]

OpenStudy (anonymous):

One side is the numerator the other the denom

OpenStudy (anonymous):

=/ the negative exponents are showing and normal sized numbers. All the negatives are exponents

OpenStudy (anonymous):

so it's something like this :\[\frac{6ab^{-3}}{6^{-1}a^{-4}b^{2}} = \frac{(6)(6)a^{4}.a}{b^3.b^2} = \frac{36a^5}{b^5}\] Coorectme if I'm wrong ^_^

OpenStudy (anonymous):

correct me*

OpenStudy (anonymous):

Oh okayyy. I understand now thanks

OpenStudy (anonymous):

np ^_^